Given the population 1,3,4,6 and 8. Suppose samples of size 3 are drawn from this population.

How many different samples can be drawn from this population?


1
Expert's answer
2022-05-17T09:26:46-0400

We have population values "1,3,4,6,8," population size "N=5" and sample size "n=2." Thus, the number of possible samples which can be drawn without replacement is

"_NC_n=\frac {N!} {(N-n)!n!}"

"_5C_3=\frac {5!} {(5-2)!2!}=\frac {120} {2!3!}=\frac {120} {12}=10"